A person is painting his house walls. He stands on a ladder with a bucket containing paint in one hand and a brush in other. Suddenly the bucket slips from his hand falls down on the floor. If the bucket with the paint had a mass of 6.0 kg and was at a height of 2.0 m at the time it slipped. How much gravitational potential energy is lost together with the paint?

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. m = 6kg, h = 2m P.E. at a height ‘2m’ = mgh = 6 × (9.8) × 2 = 117.6 J P.E. at floor = 0 Loss in P.E. = 117.6 – 0 = 117. 6 J = 118 J